# Dependency equilibria: Boundary cases and their real algebraic geometry

### v1

@@ Abstract.
This paper is a significant step forward in understanding dependency equilibria within the framework of real algebraic geometry encompassing both pure and mixed equilibria. We start by breaking down the concept for a general audience, using concrete examples to illustrate the main results. In alignment with Spohn’s original definition of dependency equilibria, we propose three alternative definitions, allowing for an algebro-geometric comprehensive study of all dependency equilibria. We give a sufficient condition for the existence of a pure dependency equilibrium and show that every Nash equilibrium lies on the Spohn variety, the algebraic model for dependency equilibria. For generic games, the set of real points of the Spohn variety is Zariski dense. Furthermore, every Nash equilibrium in this case is a dependency equilibrium. Finally, we present a detailed analysis of the geometric structure of dependency equilibria for $(2\times 2)$-games.
@@ 1. Introduction
The notion of dependency equilibria (DE) was introduced by Spohn in 2003 attempting to rationalize cooperation in Prisoner’s Dilemma. This concept suggests that players’ actions influence each other in a causal loop. This finds elucidation through reflexive decision theory, which appears as the players have a common cause in the standard non-reflexive model [13, 16]. Spohn also studied the concept [14] with the highlighting of the limited insight such as the linear loss of the mathematical study even for $(2\times 2)$-games. This started the fundamental geometric study of DE by the first author and Sturmfels in [10]. In this paper, first connection between algebraic statistics and game theory was established. The classical notion of graphical models was used to examine the so-called conditional independence equilibria that model different possible dependencies between the players. This gave rise to universality theorems for conditional independence equilibria [11, 12], similar to the universality for Nash equilibria [2]. There could be three formal objections against the general approach carried out in these publications:
@@
    The authors considered all complex points satisfying the equations for DE of a given game (that is, the complex projective Spohn variety of the game). However, the main interest of game theorists, epistemologists, and people in economics concerns, of course, the set of real non-negative points because these are the ones that potentially can be DE at all.
@@
    Spohn’s definition of DE, see [13, 14], involves conditional probabilities and, therefore, divisions by terms which are potentially equal to 00. He fixes the definition by using limits. Since the paper [10] was intended as a first step towards the geometric study of DE, this technicality was left unregarded and points where such a “division by 00” happens were not considered. Thus, only the totally mixed equilibria were studied. In particular, while it was shown that every totally mixed Nash equilibrium is a totally mixed dependency equilibrium, the generalization of this result to all Nash equilibria remained open.
@@
    The payoff tables are most of the time assumed to be generic. In real-life situations, the converse might happen.
In this paper, we work on these three points to make the results applicable for game theoretic purposes. In the next section, we present a down-to-earth explanation of our main results for the general audience with many examples.
2. Preliminaries, explanation of the results and examples
The purpose of this section is to make the results of Sections 3 and 4 and their game-theoretical relevance more accessible to an audience less familiar with algebraic geometry. It consists of explanations that intend to present the results in an informal but useful way or make them easier to read. These are accompanied by examples in which the results are demonstrated. Each explanation comes with a short description right after its numbering that tells what it is about or which result it refers to.
Explanation 2.1 (What is the Spohn variety?).
In general, a (complex algebraic) variety is, by definition, the set of all points in a finite dimensional complex vector space $\mathbb{C}^{s}$ that are the solutions of a given system of polynomial equations in $s$ indeterminates over $\mathbb{C}$. Therefore, each polynomial system gives a variety. Varieties can also be defined in projective spaces, but, also then, they look like the varieties defined above on an open dense subset.
A variety can be reducible, that is, it can be the union of two (or more) varieties strictly contained in it. In other words, the problem of solving the polynomial system translates in solving two (or more) “smaller” systems that are usually easier because they carry less information. If this cannot be done, the variety is irreducible.
A variety has a dimension which coincides with the intuitive notion of dimension: A finite point set is zero-dimensional, a curve is one-dimensional, a surface is two-dimensional, etc. As for linear systems, also for general polynomial systems, it is desirable to know the dimension of the solution set because it quantifies its size in an imaginable way.
The degree of a variety is a measure of how “wound” it is. To get the degree of a variety $V\subseteq\mathbb{C}^{s}$ of dimension $d$, ask yourself the following question:
If I want to cut $V$ with a linear variety $L$ (the solution set of a system of linear equations), what will the number of intersection points usually be if it is finite?
First, you can note that the intersection of $V$ and $L$ will usually only be finite if the dimension of $L$ is $s-d$. Second, you will see that for “most” choices of $L$ of this dimension, this number will be the same and maximal among all these numbers. It is the degree of $V$. The easiest example to consider is a curve in two-dimensional space defined by a single equation. The number of points that arise as the intersection of this curve with a line in general position gives the degree of the curve. A linear variety has degree $1$. A surface defined by a single equation of degree $m$ in $\mathbb{C}^{3}$ has degree $m$.
@@ 2. Preliminaries, explanation of the results and examples
Algebraic varieties and their generalizations (such as schemes) are well-studied in mathematics and, more specifically, in algebraic geometry, and their investigation has evolved to a far-reaching theory. For references to the topic, see, for instance, [1] introducing classical algebraic geometry and [4] developing the theory of schemes.
@@
The algebraic variety that we are concerned with in order to study DE is called the Spohn variety. We set up some notation before its definition: The ambient space of the Spohn variety is a complex projective space $\mathbb{P}^{s}$ of dimension $s=d_{1}\cdots d_{n}-1$, where $n$ is the number of players and $d_{i}$ is the number of pure strategies of player $i$. The reader not so familiar with projective spaces might just think of the ambient space as $\mathbb{C}^{s}$. Each player has a payoff table which is a tensor $X^{(i)}$ of format $d_{1}\times\cdots\times d_{n}$ with real entries. For $2$-player games, these are two $(d_{1}\times d_{2})$-matrices. Here, $X^{(i)}_{j_{1}\ldots j_{n}}$ represents the payoff that player $i$ receives when player $1$ chooses strategy $j_{1}$, player $2$ chooses strategy $j_{2}$, etc. We consider joint probability distributions over the set of pure strategies of the players: The entry $p_{j_{1}\ldots j_{n}}$ is the probability that the first player chooses the strategy $j_{1}$, the second player $j_{2}$, etc. In a DE $p$ with such entries, each player $i\in[n]=\{1,\ldots,n\}$ maximizes their conditional expected payoff conditioned on them having fixed pure strategy $k\in[d_{i}]$. In the following, the hat means that the $i$th sum is removed:
@@
Let $X$ be a $(d_{1}\times\cdots\times d_{n})$-game in normal form. A joint probability distribution $p$ is called a dependency equilibrium (DE) if and only if $\mathbb{E}^{(i)}_{k}(p)\geq\mathbb{E}^{(i)}_{k^{\prime}}(p)$ for all $k,k^{\prime}\in[d_{i}]$ and for all $i\in[n]$.
One may describe the totally mixed dependency equilibria of a game $X$ as the points of an algebraic variety associated to $X$, the Spohn variety $\mathcal{V}_{X}$ of $X$, that have positive real coordinates. The restriction to totally mixed dependency equilibria was needed as the denominators of the conditional expected payoffs might vanish. We will see later (Corollary 3.11) that the Spohn variety also plays a prominent role in the study of general dependency equilibria.
@@
The Spohn variety $\mathcal{V}_{X}$ of the game $X$ is defined as the $2\times 2$ minors of the matrices following $d_{i}\times 2$ matrices of linear forms $M_{1},\ldots,M_{n}$:
@@
In [10, Theorem 6], it is shown that for games in a “large” set (in particular, for games in an open dense subset of the set of all games, where a normal form game is considered as a point given by its payoff tables), the Spohn variety is irreducible of codimension $d_{1}+\ldots+d_{n}-n$ (that is, of dimension $s$ minus this number) and degree $d_{1}\cdots d_{n}$. For games outside this set the dimension is at least the first number and the degree is at most the second number. This follows from Bézout’s Theorem, a standard result in algebraic geometry. Thus, understanding the Spohn variety is the first step in understanding how the set of dependency equilibria of a game might look like.
@@
Let us consider $(2\times 2)$-games in normal-forms, that is, games with two players each of which has two pure strategies, also called bimatrix games. For simplicity, we denote the players’ $2\times 2$ payoff matrices as
@@
which are the two following equations
@@
    An example of a game that is “generic” in the sense that the Spohn variety is an irreducible curve of degree $4$ is given by game 114 in the Periodic Table of the $2\times 2$ Games of Goforth and Robinson, see [3]:
@@
    Its real part $\mathcal{V}(\mathbb{R})$ intersects the tetrahedron $\overline{\Delta}$ of all strategies, as shown in Figure 1. The first irreducible component enters $\overline{\Delta}$ at the pure strategy $(1,0,0,0)$ and leaves it at $(0,0,0,1)$. The second irreducible component intersects $\overline{\Delta}$ at the other two pure strategies but enters it only later at two mixed strategies of the form $(p_{11},0,p_{12},p_{22})$ and $(p_{11},p_{21},0,p_{22})$. We will argue later in Corollary 3.11 that for generic games, including this game, the set of DE is given exactly by the common points of $\mathcal{V}(\mathbb{R})$ and $\overline{\Delta}$.
@@
Explanation 2.5 (concerning Theorem 3.2 and Corollary 3.4).
A game theorist would probably interpose in view of [10, Theorem 6]: I want to understand totally mixed dependency equilibria. These are at most the real points of the Spohn variety. So, what can you tell me about the real points, their dimension and degree? Are they an irreducible set?
@@
In general, the set of points with real coordinates of a variety can have very strange properties. It might be even empty. However, the “in particular” of Theorem 3.2 tells us that for generic games $X$ (that is, again, for an open dense subset of the set of all games) the real part of the Spohn variety $\mathcal{V}_{X}$ is as large and looks structurally exactly as one would expect from looking at $\mathcal{V}_{X}$. Corollary 3.4 then uses this statement to infer for generic games that the set of real points on $\mathcal{V}_{X}$ is also irreducible, it has (real) dimension $s-(d_{1}+\ldots+d_{n}-n)$ and degree $d_{1}\cdots d_{n}$..
Explanation 2.6 (Three types of DE, Definition 3.6).
The notion of DE that we propose in Definition 3.6(1) keeps the conceptional game theoretic logic of Spohn’s original definition. It broadens the notion naturally in a purely mathematical manner by allowing, in order to approximate DE, instead of positive real points, as proposed by Spohn, arbitrary points for which the rational expressions $\mathbb{E}^{(i)}_{k}(p)$ defining DE are sensible. (2) Analytic DE is a notion that is intended to simplify the study of DE. (3) geometric DE are their algebro-geometric analogue. All three concepts turn out to be sensible in view of Corollary 3.11.
Explanation 2.7 (Bounds for the set of DE, Corollary 3.11).
@@
Corollary 3.11 shows that an investigation of DE from the perspective of algebraic geometry is indeed sensible. Every DE is a positive real point on the Spohn variety $\mathcal{V}=\mathcal{V}_{X}$ and every strategy profile that lies on an irreducible component of the Spohn variety that contains a point, for which the rational expression $\mathbb{E}^{(i)}_{k}(p)$ of DE are well-defined, is a DE. The union of all such irreducible components is written as $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}$ in Corollary 3.11 where $\mathcal{W}\subseteq\mathbb{P}^{d_{1}\times\ldots\times d_{n}-1}$ is the union of hyperplanes defined by the polynomial
$s=\prod_{i,k}p_{+\ldots+k+\ldots+}.$
These bounds are computationally very helpful because they can, given enough time, be computed by a computer algebra program such as Macaulay2. An explicit example of code for $(2\times 2)$-games is shown in Example 2.8 below.
We will see in Section 4 that none of the two bounds does necessarily agree with the set of DE. However, they do agree for generic games, which means for games outside of some algebraic variety of games. Here, we view the set of all $(d_{1}\times\cdots\times d_{n})$-games in normal form as the real vector space $\mathbb{R}^{n\times d_{1}\times\cdots\times d_{n}}$. It would be very interesting to explicitly determine this variety for games of a given size. We do this for $(2\times 2)$-games in Theorem 4.9, where this exceptional variety turns out to just be a union of hyperplanes in $\mathbb{R}^{8}$.
@@
The following is a code for computations with the computer algebra software Macaulay2 [5] of lower and upper bounds for the set of DE of a $(2\times 2)$-game.

### v3

@@ Abstract.
This paper is a significant step forward in understanding dependency equilibria within the framework of real algebraic geometry encompassing both pure and mixed equilibria. In alignment with Spohn’s original definition of dependency equilibria, we propose two alternative definitions, allowing for an algebro-geometric comprehensive study of all dependency equilibria. We give a sufficient condition for the existence of a pure dependency equilibrium and show that every Nash equilibrium lies on the Spohn variety, the algebraic model for dependency equilibria. For generic games, the set of real points of the Spohn variety is Zariski dense. Furthermore, every Nash equilibrium in this case is a dependency equilibrium. Finally, we present a detailed analysis of the geometric structure of dependency equilibria for $(2\times 2)$-games.
@@ 1. Introduction
A well-known concept in game theory, the Nash equilibrium, describes a state in which no player can improve their payoff by unilaterally changing their strategy. However, in some games, Nash equilibria do not lead to the most optimal outcomes, as in the case of the Prisoner’s Dilemma. Spohn addresses this issue by introducing the notion of dependency equilibria (DE) in 2003 to rationalize cooperation in the Prisoner’s Dilemma. This concept suggests that players’ actions influence each other in a causal loop. The notion of DE can be understood through reflexive decision theory, where players appear to share a common cause in the standard non-reflexive model [13, 16]. Spohn studied DE [14], highlighting its non-trivial and fundamentally different nature compared to Nash equilibria, as even in $(2\times 2)$-games, it is no longer described by linear polynomials [8]. This started the fundamental geometric study of DE by the first author and Sturmfels in [10]. In this paper, the first connection between algebraic statistics and game theory was established. The classical notion of graphical models was used to examine the so-called conditional independence equilibria that model different possible dependencies between the players. This gave rise to universality theorems for conditional independence equilibria [12, 11], similar to the universality for Nash equilibria [2]. In the following, we present three possible formal objections against the general approach carried out in these publications and how we address them throughout the paper to make the results applicable for game theoretic purposes.
@@
    The authors considered all complex points satisfying the equations for the dependency equilibria of a given game, that is, the complex projective Spohn variety of the game ([10, Theorem 3.2], Definition 2.2). However, the main interest of game theorists, epistemologists, and people in economics concerns, of course, the set of real non-negative points because these are the ones that potentially can be DE at all.
    1. $\bullet$
      In Theorem 3.2, we prove that for generic games, the set of real points of the Spohn variety is Zariski dense. Consequently, the results on the complex Spohn variety from [10, Theorem 3.2] for generic games also hold for its set of real points.
@@ 1. $\bullet$
    Spohn’s definition of DE, see [13, 14], involves conditional probabilities and, therefore, divisions by terms which are potentially equal to 00. He fixes the definition by using limits. Since the paper [10] was intended as a first step towards the geometric study of DE, this technicality was left unregarded and points where such a “division by 00” happens were not considered. Thus, only the totally mixed equilibria were studied.
    1. $\bullet$
      In Definition 3.6, we give two definitions for general DE and we prove that every dependency equilibrium of a game lies on the Spohn variety (Lemma 3.9). Theorem 3.22 gives a criterion, in terms of linear algebra, for the existence of a whole family of totally mixed DE around a pure strategy. This criterion also serves as a sufficient condition for the existence of a pure DE.
    2. $\bullet$
      In Theorem 3.19, we prove that every Nash equilibrium of a game $X$ lies on the Spohn variety of $X$. Moreover, if no irreducible component of the Spohn variety is contained in the hyperplanes generated by the denominators appearing in the definition of DE (conditional probabilities that are potentially zero), then every Nash equilibrium is a dependency equilibrium (Corollary 3.20).
@@ 2. $\bullet$
    The payoff tables are most of the time assumed to be generic. In real-life situations, many common models (e.g. Example 2.3(b), Example 3.5) are not generic.
    1. $\bullet$
      In addition to the previous two points, we show that the Spohn variety can include points on the boundary of the probability simplex that are not DE. Moreover, there may be more DE than just the points on the Spohn variety whose conditional payoffs are well-defined (Theorem 3.10). However, generically, these distinctions do not need to be considered (Proposition 3.16). Throughout the paper, we also present detailed computations of various non-generic examples of games. In Section 4, we specifically examine arbitrary $(2\times 2)$-games.
2. Preliminaries and first examples
@@
The purpose of this section is to introduce some game-theoretic notation for the following sections and its interpretation in terms of algebraic varieties. These geometric objects and their generalizations (such as schemes) are well-studied in mathematics and, more specifically, in algebraic geometry, and their investigation has evolved to a far-reaching theory. For references to the topic, see, for instance, [1] introducing classical algebraic geometry and [4] developing the theory of schemes. The algebraic variety that we are concerned with in order to study dependency equilibria (DE) is called the Spohn variety. We set up some notation before its definition.
@@
The ambient space of the Spohn variety is a complex projective space $\mathbb{P}^{s}$ of dimension $s=d_{1}\cdots d_{n}-1$, where $n$ is the number of players and $d_{i}$ is the number of pure strategies of player $i$. The reader not so familiar with projective spaces might just think of the ambient space as $\mathbb{C}^{s}$. Each player has a payoff table which is a tensor $X^{(i)}$ of format $d_{1}\times\cdots\times d_{n}$ with real entries. For $2$-player games, these are two $(d_{1}\times d_{2})$-matrices. Here, $X^{(i)}_{j_{1}\ldots j_{n}}$ represents the payoff that player $i$ receives when player $1$ chooses strategy $j_{1}$, player $2$ chooses strategy $j_{2}$, etc. The tuple $X=(X^{(i)})$ is typically called a game in normal form and it can be interpreted as a point in $\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$.
We consider (joint) strategies $p$, that is, joint probability distributions over the set of pure strategies of the players: The entry $p_{j_{1}\ldots j_{n}}$ is the probability that the first player chooses the strategy $j_{1}$, the second player $j_{2}$, etc. We will frequently interpret $p$ as a point in $\mathbb{P}^{s}$ that admits a representation by non-negative real coordinates. We denote by $\overline{\Delta}$ the set of all joint strategies and by $\Delta$ the set of all totally mixed strategies, that is, strategies with all entries $p_{j_{1}\ldots j_{n}}>0$. The finitely many strategies with one entry equal to $1$ (and hence all others equal to 00) are called pure. A DE is a strategy $p$ where each player $i\in[n]=\{1,\ldots,n\}$ maximizes their conditional expected payoff $\mathbb{E}^{(i)}_{k}(p)$ conditioned on them having fixed pure strategy $k\in[d_{i}]$. In the following, the hat means that the $i$th sum is removed:
@@
Let $X$ be a $(d_{1}\times\cdots\times d_{n})$-game in normal form. A joint probability distribution $p$ is called a dependency equilibrium (DE) if $\mathbb{E}^{(i)}_{k}(p)\geq\mathbb{E}^{(i)}_{k^{\prime}}(p)$ for all $k,k^{\prime}\in[d_{i}]$ and for all $i\in[n]$.
The above definition uses the condition $\mathbb{E}^{(i)}_{k}(p)\geq\mathbb{E}^{(i)}_{k^{\prime}}(p)$ for all $k,k^{\prime}\in[d_{i}]$. Hence, similarly to [10, Equation (4)], we could also replace the inequality by an equality.
One may describe the totally mixed DE of a game $X$ as the points of an algebraic variety associated to $X$, the Spohn variety $\mathcal{V}_{X}$ of $X$, that have positive real coordinates [10, Theorem 3.2]. The restriction to totally mixed DE was needed as the denominators of the conditional expected payoffs might vanish. We will treat the general case of not necessarily totally mixed strategies in Section 3 and we will see that the Spohn variety also plays a prominent role there.
@@
The Spohn variety $\mathcal{V}_{X}\subseteq\mathbb{P}^{s}$ of the game $X$ is the complex projective variety defined by the $(2\times 2)$-minors of the following $(d_{i}\times 2)$-matrices $M_{1},\ldots,M_{n}$ of linear forms:
@@
Throughout the paper, we will use the term “generic” in the sense of algebraic geometry. More precisely, a property or a statement holds generically if it holds on an open (and hence dense) subset of the ambient space with respect to the Zariski topology. In our case, the ambient space will always be the real affine space $\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$ of all games of a fixed size. In this setting, a statement like “property $P$ holds for generic games” translates as “the set of games $X\in\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$ satisfying property $P$ is Zariski open”. Note that the usual notion of a generic game from game theory is similar to our notion of genericity, where $P$ is defined as “the set of Nash equilibria of $X$ consists of isolated points”. However, these two concepts of genericity may differ and should be studied separately. One example is “Bach or Stravinsky” (Example 3.5): Whereas this game has isolated points as its set of Nash equilibria, it is not generic in the sense of DE, since its Spohn variety is a reducible curve.
In [10, Theorem 3.2], it is shown that, for generic games, the Spohn variety is irreducible of codimension $d_{1}+\ldots+d_{n}-n$ and degree $d_{1}\cdots d_{n}$ in $\mathbb{P}^{s}$. In any case, the dimension is at least the first number and the degree is at most the second number. This follows from Bézout’s Theorem, a standard result in algebraic geometry. Understanding the Spohn variety is the first step in understanding how the set of DE of a game might look like.
@@
Let us consider $(2\times 2)$-games in normal form, that is, games with two players each of which has two pure strategies, also called bimatrix games. For simplicity, we denote the players’ $(2\times 2)$-payoff matrices as
@@
which give the two following defining equations
@@
    An example of a game that is generic in the sense that the Spohn variety is an irreducible curve of degree $4$ is given by game 114 in the Periodic Table of the $2\times 2$ Games of Goforth and Robinson, see [3]:
@@
    Its real part $\mathcal{V}(\mathbb{R})$ intersects the tetrahedron $\overline{\Delta}$ of all strategies, as shown in Figure 1. The first irreducible component enters $\overline{\Delta}$ at the pure strategy $(1,0,0,0)$ and leaves it at $(0,0,0,1)$. The second irreducible component intersects $\overline{\Delta}$ at the other two pure strategies but enters it only later at two mixed strategies of the form $(p_{11},0,p_{12},p_{22})$ and $(p_{11},p_{21},0,p_{22})$. We will argue later in Theorem 3.10 that for generic games, including this game, the set of DE is given exactly by the common points of $\mathcal{V}(\mathbb{R})$ and $\overline{\Delta}$.
@@
3. Results for arbitrary games
Real points of the Spohn variety
In the following, we show that the Spohn variety $\mathcal{V}=\mathcal{V}_{X}$ of a game $X$, that is, the variety of all complex projective points satisfying the equations of a dependency equilibrium, always has smooth points with real coordinates. This enables us to apply the following fact which, in turn, makes it possible to study the set of all DE from the point of view of complex projective varieties. As a consequence, we can apply the results of [10] to the case that is actually relevant for game-theoretical applications, namely, to real points on the Spohn variety.
Fact 3.1.
[9, Theorem 2.2.9] Let $V$ be an irreducible complex affine variety defined by real polynomials. If $V$ contains a smooth point with real coordinates then $V(\mathbb{R})$, the set of real points of $V$, lies Zariski dense in $V$.
@@ 3. Results for arbitrary games
We can apply this by taking a dense open affine subset of the ambient projective space and thereby assuming that $V=\mathcal{V}$ is affine. In general, the set of points with real coordinates of a variety can have very strange properties and might even be empty. However, the “in particular” of Theorem 3.2 tells us that for generic games $X$ the real part of the Spohn variety $\mathcal{V}_{X}$ is as large and looks structurally exactly as one would expect from looking at $\mathcal{V}_{X}$. Corollary 3.4 then uses this statement to infer that the set of real points on $\mathcal{V}_{X}$ is generically irreducible, it has (real) dimension $s-(d_{1}+\ldots+d_{n}-n)$ and degree $d_{1}\cdots d_{n}$.
Theorem 3.2.
Proof.
For $(2\times 2)$-games, this is just Proposition 4.1.
In all other cases, we use the vector bundle parametrization $\phi:\mathcal{V}_{X}\dashrightarrow(\mathbb{P}^{1})^{n}$ appearing in the proof of [10, Theorem 3.4] and can write
Regarding game-theoretic applications, it might be interesting to find smooth real points on every irreducible component of the Spohn variety of a game, as this would prove the Zariski density of real points of the Spohn variety for arbitrary games. Consequently, the Spohn variety would, in general, structurally reflect DE. In particular, this has been proven for $(2\times 2)$-games in [8, Lemma 3.5].
Question 3.3.
The following is the combination of Theorem 3.2 with [10, Theorem 3.2].
Corollary 3.4.
Passing to boundary cases
Spohn [13, 14] defines DE on the boundary of $\overline{\Delta}$ by using limits. The definition we use is slightly more restrictive but coincides with Spohn’s for generic games. A DE is a point $p=(p_{j_{1}\ldots j_{n}})$ in $\overline{\Delta}$ that is the limit of a sequence $(p^{(r)})_{r\in\mathbb{N}}$ in $\Delta$ satisfying
for all $i\in\{1,\ldots,n\}$ and $k,k^{\prime}\in\{1,\ldots,d_{i}\}$. In the special case of $(2\times 2)$-games with the notation of [10, Example 2.2], these conditions read as follows:
Figure 2. For “Bach or Stravinski”, the intersection (red) of the real part of the Spohn variety $\mathcal{V}$ (blue) with $\overline{\Delta}$ (orange) consists of two isolated points and two line segments (red). The isolated points are Nash equilibria but they are not DE in our sense, see Example 3.5.
Example 3.5 (Bach or Stravinski).
We use the payoff tables $(2,0,0,1)$ and $(1,0,0,2)$ as in [14, Section 3], see also Figure 2. The equations (1) translate as
The two Nash equilibria $(1,0,0,0)$ and $(0,0,0,1)$ cannot be approximated by points in $\Delta$ satisfying the equations above. Indeed, assume to the contrary that the sequence $p^{(r)}$ converges to $(1,0,0,0)$ and the above equations are satisfied. Then, $\lim_{r\to\infty}p_{11}^{(r)}=1$ and all other $p_{ij}^{(r)}$ converge to 00. This means that $\lim_{r\to\infty}\frac{2p^{(r)}_{11}}{p^{(r)}_{11}+p^{(r)}_{12}}=2=\lim_{r\to\infty}\frac{p^{(r)}_{22}}{p^{(r)}_{21}+p^{(r)}_{22}}$. Since all the $p_{ij}^{(r)}$ are strictly positive, it follows that $\frac{1}{2}=\lim_{r\to\infty}\frac{p^{(r)}_{21}+p^{(r)}_{22}}{p^{(r)}_{22}}=\lim_{r\to\infty}\frac{p^{(r)}_{21}}{p^{(r)}_{22}}+1$. So $\lim_{r\to\infty}\frac{p^{(r)}_{21}}{p^{(r)}_{22}}=-\frac{1}{2}$ and hence this last sequence must get negative for $r$ sufficiently large. This is a contradiction to $p_{ij}^{(r)}>0$.
We will now slightly generalize DE by also allowing limits from outside $\overline{\Delta}$. As a consequence, Nash equilibria will generically be DE. Then, we will describe the connection to the algebraic setting of the Spohn variety. Let $X$ be a game given by its payoff tensors $(X^{(i)}_{j_{1}\ldots j_{n}})$, $\mathcal{V}=\mathcal{V}_{X}\subseteq\mathbb{P}^{d_{1}\times\ldots\times d_{n}-1}$ the Spohn variety of $X$. Moreover, let $\mathcal{W}\subseteq\mathbb{P}^{d_{1}\times\ldots\times d_{n}-1}$ be the union of hyperplanes defined by the polynomial
$s=\prod_{i,k}p_{+\ldots+k+\ldots+},$
where $i$ is the position in which $k$ appears. Note, that $\mathcal{W}$ is exactly the union of those hyperplanes outside which all the conditional expected payoffs $\mathbb{E}_{k}^{(i)}(p)$ are well-defined. For a subset $M\subseteq\mathbb{C}^{d_{1}\times\ldots\times d_{n}}$ (respectively $\mathbb{P}^{d_{1}\times\ldots\times d_{n}-1}$ or $\mathbb{R}^{d_{1}\times\ldots\times d_{n}}$), we denote by $\overline{M}$ its usual euclidean closure, that is, the set of all points that are limits of sequences in $M$, and by $\overline{M}^{\text{Zar}}$ its Zariski closure, that is, the set of all points that are common zeros of all polynomials vanishing on $M$. It is a basic fact that $\overline{M}\subseteq\overline{M}^{\text{Zar}}$ because every algebraic set is closed in the euclidean topology.
Definition 3.6.
  1. (1)
  2. (2)
    An algebraic dependency equilibrium (aDE) is a point in $\overline{\Delta}\cap\overline{(\mathcal{V}\setminus\mathcal{W})}^{\text{Zar}}$.
For all $(2\times 2)$-games in [14, Section 3] (Matching pennies, Bach or Stravinski, Hawk and Dove, Prisoner’s Dilemma), it can be easily checked that DE and aDE coincide. For generic games, we can replace complex tensors by real tensors in Definition 3.6 and still deduce results analogous to the contents of this section, see Remark 3.13.
Remark 3.7.
  1. (1)
    We could equivalently define aDE as points in $\overline{\Delta}\cap\overline{\mathcal{V}\setminus\mathcal{W}}$. This follows from the general fact that the Zariski closure and the Euclidean closure coincide for an open subset of a variety, see [7, Exposé XII, Prop. 2.2].
  2. (2)
    It follows from (1) that every aDE is a DE, because it can be approximated by a sequence in $\mathcal{V}\setminus\mathcal{W}$.
Remark 3.8.
Lemma 3.9.
Proof.
The following is now a direct consequence of Remark 3.7(2) and Lemma 3.9.
Theorem 3.10.
For the set $\text{DE}(X)$ of DE of a game $X$, there are inclusions
$\overline{\Delta}\cap\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}=\overline{\Delta}\cap\overline{\mathcal{V}\setminus\mathcal{W}}\subseteq\text{DE}(X)\subseteq\overline{\Delta}\cap\mathcal{V},$
where $\mathcal{V}=\mathcal{V}_{X}$ is the Spohn variety of $X$.
@@
Theorem 3.10 shows that an investigation of DE from the perspective of algebraic geometry and computational commutative algebra is indeed sensible. Every DE is a positive real point on the Spohn variety $\mathcal{V}=\mathcal{V}_{X}$ and every strategy that lies on an irreducible component of the Spohn variety that contains a point, for which the rational expression $\mathbb{E}^{(i)}_{k}(p)$ of DE are well-defined, is a DE. The union of all such irreducible components is written as $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}$ in Theorem 3.10. These bounds are very helpful from a computational perspective because they can, given enough time, be computed by a computer algebra program such as Macaulay2. An explicit sample of code for $(2\times 2)$-games is shown in Example 3.11 below.
We will see in Example 3.12 and Section 4 that neither of the two bounds of Theorem 3.10 does necessarily agree with the set of DE. However, they do agree for generic games as we will show in Proposition 3.16.
@@
The following is code for computations with the computer algebra software Macaulay2 [5] of lower and upper bounds for the set of DE of a $(2\times 2)$-game.